OpenStudy (anonymous):

Bill had 25 cars, Sylvia wanted 90% of his cars! How many cars does Bill have left over!

OpenStudy (zarkon):

0...because Sylvia decided that 90% was just not enough and wanted them all

OpenStudy (anonymous):

lol

OpenStudy (diyadiya):

25x90/100=90/4=22.5

OpenStudy (anonymous):

lol i just made this to keep you guys busy lol :)

OpenStudy (zarkon):

I feel so used...just like Bill

OpenStudy (anonymous):